# NUDT4MSTAR: A New Dataset and Benchmark Towards SAR Target Recognition in the Wild

### v1

@@ 4. $\bullet$
In this section, we describe the dataset construct program and then provide the details on NUDT4MSTAR in terms of key aspects, including data acquisition, professional system, statistical analysis, and dataset value.
@@ 3.1 Construct program
Target. According to the Chinese and European vehicle classification standards [75, 17], First, we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each category into various target classes based on size and special structure and collect as different types as possible. Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types. It covers the most of civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets in Fig. 5.
@@
Table II shows the open-source data list. After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction. We commissioned the Institute of Stealthy Engineering Technology Company in Beijing to complete the data collection and annotation.
@@ 3.3 Professional system
Figure 8: Distribution of object size. Our dataset has a richer object sizes compared to another civilian vehicle dataset (Gotcha). It is clear that the size and ratio of the vehicles are within a certain range due to real-world constraints, such as official regulations and driving requirements, and we will continue to expand the object categories beyond vehicles.
@@ 3.4 Statistical Analysis
Figure 13: Complex data. Here is a data example of Vehicle 5 (Great Wall Voleex C50).
@@
Band and polarization. Table I shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings. However, Fig. 12 demonstrates that this setting is not the only one to consider, as band and polarization have a noticeable effect on target scattering. We provide two bands and quad polarization to support discussing these sensor parameters.
Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format as Fig. 13 shown.
@@
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors carried out the data annotation over about four months. The entire project cost about CNY 1.45 million.
@@ 3.5 Dataset value
Robustness recognition. Our data has diverse acquisition conditions, and each image is labeled with specific information. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene, and sensor) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [99].
@@ 4 Benchmarks
TABLE III: SOC and EOC settings of NUDT4MSTAR. The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition. Simple scenes are sandstone and bare soil, and complex scenes are urban, factory, and woodland. The separate labeling of the ground distance and slant distance images results in their numbers of annotations not being strictly corresponding. # Types.: Number of object types. Dep.: Depression angle. Azi.: Target azimuth angle. Pol.: Polarization. # Img. (Ground): Number of images in ground range. # Img. (Slant): Number of images in slant range.
@@
We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table III. These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts. The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
4.1 Experimental settings
Here, we describe the experimental settings in Table III, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
@@ 4.1 Experimental settings
Transfer learning. To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table IV, and the weights are used for other SAR target classification datasets as initialization for fine-tuning. Fine-tuning parameter settings follow our previous paper [48], and the baseline is the weight of ImageNet supervised pre-training for these models.
@@ 4.2 Results and analyses
TABLE IV: Classification results. We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples. Gray color methods use complex images, and magnitude image methods are sorted by network architecture, such as CNN and Transformer. Bolded text indicates the best result, while underlined text is the next best result.
@@
Classification. From different methods in Table IV, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40. Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene. The model recognition performance and robustness of the latest 2020s CNN architecture ConvNeXt are superior to the previous CNN models. ConvNeXt is more robust under EOC-Deprssion than other methods, and ViT is insensitive to azimuthal angle change. Our previously proposed method, HDANet, improves the attention mechanism and uses a segmentation task so it can be used in different scenarios. In addition, our proposed LDSF introduces a scattering graph structure that effectively improves the discrimination of targets and clutter. SARATR-X is a pre-trained self-supervised learning model on SAR with the HiViT architecture, and its performance is significantly improved compared to HiViT. Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
The different experimental settings in Table IV show that there are still great challenges in SAR target recognition. As the number of fine-grained classes increased, previous methods were no longer able to achieve nearly 99% accuracy on MSTAR’s SOC. SOC-50 is closer to the actual situation, as we are unlikely to obtain a large number of target samples under different conditions. The test set accuracy tends to decrease linearly as the training samples are reduced on a logarithmic scale in Fig 14. In the difficult scene of EOC-Scene, deep learning encounters great challenges, especially in the background clutter interference inside the trees inside the woodland. The significant deterioration in accuracy indicates that deep learning models trained on simple backgrounds have difficulty distinguishing between targets and complex background clutter. The use of complex information or the inclusion of auxiliary tasks such as segmentation during training can enhance the recognition ability, but the alteration of target features due to background clutter interference is still a difficult challenge. We observe a significant decrease using the complex data compared to the magnitude data under angles and sensor parameter variations. These results are because the complex phase is more sensitive to these parameters’ transformations, and the magnitude of the data is additionally preprocessed to enhance the image’s visual quality. Changes in imaging angle have a greater impact on target characteristics than band and polarization, especially large depression and azimuth angle changes. Different imaging angles lead to variations in target and background clutter, as well as different occlusions and scattering resulting from changes in imaging geometry. We have not found a good method to deal with the above scene and angle changes. The shift in target features due to band and polarization can be solved very effectively by ConvNeXt and SARATR-X. However, due to cost issues, we have not discussed the effect of different platform resolutions on recognition in the current version, which is another important issue.
TABLE V: Detection results. We use mAP50 [55] as a metric, i.e., the average precision at IoU 0.5 because the horizontal bounding box does not exactly fit the contours of the different targets. Bolded text indicates the best result, while underlined text is the next best result.
@@
Detection. As illustrated in Table V, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions. In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions. However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions. In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios. This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency. The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
@@
Transfer learning. Fig 15 shows how our dataset can help different models benefit from feature learning and be used for target recognition in other datasets. We use Table’s SOC-40 trained model weights as initializations on other target datasets with 5-shot sample experiments [48]. The fine-tuning results show that similar ground target-type datasets, such as vehicles and airplanes, would greatly benefit from training on our data. However, the transfer effect is not obvious for the ship target class dataset in the sea scene.
@@ 4.3 Discussion
Robustness under different conditions. As Table IV and  V show, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
@@
Incorporation with multi-dimensional information. Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table IV. Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.

### v2

@@ 3.1 Construct program
Target. First of all, according to the Chinese and European vehicle classification standards [82, 19], we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each category into various target classes based on size and special structure and collect as many different types as possible. Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types. It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
@@
Table III shows the open-source data list. After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction. We commissioned the Institute of Stealthy Engineering Technology Company in Beijing to complete the data collection and annotation.
@@ 3.4 Statistical Analysis
Figure 8: Distribution of object size. Our dataset has a richer object size distribution compared to another civilian vehicle dataset (Gotcha). It is clear that the size and ratio of the vehicles are within a certain range due to real-world constraints, such as official regulations and driving requirements, and we will continue to expand the object categories beyond vehicles.
@@
Band and polarization. Table II shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings. However, Fig. 12 demonstrates that this setting is not the only one to consider, as band and polarization have a noticeable effect on target scattering. We provide two bands and quad polarization to support discussing these sensor parameters.
Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format.
@@
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors perform the data annotation about four months.
@@ 3.5 Dataset value
Robustness recognition. Our data has diverse acquisition conditions, and each image is labeled with specific information. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene and sensor sensors) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [109].
@@ 4 Benchmarks
We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table IV. These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts. The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
4.1 Experimental settings
Here, we describe the experimental settings in Table IV, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
TABLE IV: SOC and EOC settings of NUDT4MSTAR. The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition. Simple scenes are sandstone and bare soil, and complex scenes are urban, factory, and woodland. The separate labeling of the ground distance and slant distance images results in their numbers of annotations not being strictly corresponding. # Types.: Number of object types. Dep.: Depression angle. Azi.: Target azimuth angle. Pol.: Polarization. # Img. (Ground): Number of images in ground range. # Img. (Slant): Number of images in slant range.
@@ 4.1 Experimental settings
Transfer learning. To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table V, and the weights are used for other SAR target classification datasets as initialization for fine-tuning. Fine-tuning parameter settings follow our previous paper [53], and the baseline is the weight of ImageNet supervised pre-training for these models.
@@ 4.2 Results and analyses
TABLE V: Classification results. We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples. Gray color methods use complex images, and magnitude image methods are sorted by network architecture, such as CNN and Transformer. Bolded text indicates the best result, while underlined text is the next best result.
@@
Classification. From different methods in Table V, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40. Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene. The model recognition performance and robustness of the latest 2020s CNN architecture ConvNeXt are superior to the previous CNN models. ConvNeXt is more robust under EOC-Deprssion than other methods, and ViT is insensitive to azimuthal angle change. Our previously proposed method, HDANet, improves the attention mechanism and uses a segmentation task so it can be used in different scenarios. In addition, our proposed LDSF introduces a scattering graph structure that effectively improves the discrimination of targets and clutter. SARATR-X is a pre-trained self-supervised learning model on SAR with the HiViT architecture, and its performance is significantly improved compared to HiViT. Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
The different experimental settings in Table V show that there are still great challenges in SAR target recognition. As the number of fine-grained classes increased, previous methods were no longer able to achieve nearly 99% accuracy on MSTAR’s SOC. SOC-50 is closer to the actual situation, as we are unlikely to obtain a large number of target samples under different conditions. The test set accuracy tends to decrease linearly as the training samples are reduced on a logarithmic scale in Fig 13. In the difficult scene of EOC-Scene, deep learning encounters great challenges, especially in the background clutter interference inside the trees inside the woodland. The significant deterioration in accuracy indicates that deep learning models trained on simple backgrounds have difficulty distinguishing between targets and complex background clutter. The use of complex information or the inclusion of auxiliary tasks such as segmentation during training can enhance the recognition ability, but the alteration of target features due to background clutter interference is still a difficult challenge. We observe a significant decrease using the complex data compared to the magnitude data under angles and sensor parameter variations. These results are because the complex phase is more sensitive to these parameters’ transformations, and the magnitude of the data is additionally preprocessed to enhance the image’s visual quality. Changes in imaging angle have a greater impact on target characteristics than band and polarization, especially large depression and azimuth angle changes. Different imaging angles lead to variations in target and background clutter, as well as different occlusions and scattering resulting from changes in imaging geometry. We have not found a good method to deal with the above scene and angle changes. The shift in target features due to band and polarization can be solved very effectively by ConvNeXt and SARATR-X. However, due to cost issues, we have not discussed the effect of different platform resolutions on recognition in the current version, which is another important issue.
@@
Detection. As illustrated in Table VI, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions. In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions. However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions. In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios. This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency. The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
TABLE VI: Detection results. We use mAP50 [61] as a metric, i.e., the average precision at IoU 0.5 because the horizontal bounding box does not exactly fit the contours of the different targets. Bolded text indicates the best result, while underlined text is the next best result.
@@
Transfer learning. Fig 14 shows how our dataset can help different models benefit from feature learning and be used for target recognition in other datasets. We use Table’s SOC-40 trained model weights as initializations on other target datasets with 5-shot sample experiments [53]. The fine-tuning results show that similar ground target-type datasets, such as vehicles and airplanes, would greatly benefit from training on our data. However, the transfer effect is not obvious for the ship target class dataset in the sea scene. In recent years, the number of emerging satellite ship datasets has been much larger than that of other ground targets. This data can greatly compensate for the lack of high-resolution ground target samples in the existing SAR target datasets and advance the construction of an SAR ATR foundation model.
@@ 4.3 Discussion
Robustness under different conditions. As Table V and  VI show, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
@@
Incorporation with multi-dimensional information. Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table V. Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.